# Photon distillation schemes with reduced resource costs based on multiphoton Fourier interference

### v3

@@ I Introduction
The key resource for all linear-optical quantum computing schemes is photon indistinguishability [8, 9, 10], i.e. the degree to which the photons in the quantum computer are identical in the degrees of freedom of those photons which are not being used to encode information. Photonic indistinguishability is essential because while linearity forbids an explicit photon-photon interaction, it allows for multi-particle quantum interference, which is sufficient for universal quantum computation when combined with measurements and fast feedback. The archetypal example of such interference is the Hong-Ou-Mandel effect [11]. Since photon indistinguishability governs the degree to which multi-particle interference occurs, it is the crucial resource for photonic quantum computing in its modern implementations [12, 13, 14, 15], and imperfect indistinguishability gives rise to logical errors in the computation.
@@
In this work, we present a photon distillation scheme with significantly reduced resource costs, rendering the protocol feasible for implementation in photonic quantum computers. Our scheme is based on multiphoton interference in Fourier matrices. In contrast to the state of the art [17], our scheme can achieve arbitrary reduction of errors in a single round of the protocol, thereby eliminating the need for concatenation of multiple distillation steps or intermediate fast feedback. Up to a conjecture in the theory of matrix permanents, we show that the resource costs as expressed in the overall number of photons required for the protocol are linear in the required error reduction, and the success probability of the protocol tends to a constant as more photons are used. These two facts render our protocol highly scaleable: to achieve an asymptotic reduction in indistinguishability error by at least a factor of 100 required 59,000 photons with the state-of-the-art scheme, whereas our new scheme only requires 400 photons. We show that our scheme is asymptotically optimal in the degree of error reduction per photon consumed.
@@ II Interference of partially distinguishable photons
We begin by introducing a model for partially distinguishable photons. The orthogonal bad-bit model describes single photon states produced by imperfect sources [16]. The internal error modes do not contribute to multiphoton interference because of their mutual orthogonality, thus allowing us to expand the imperfect single photon states $\rho(\epsilon)$ into equivalent sums of fully indistinguishable $\rho_{\text{i}}$ and fully distinguishable $\rho_{\text{d}}$ photons [24]
@@
where $\epsilon$ is the parameter used in this work for the partial distinguishability error  [^1] The orthogonal bad-bit model is equivalent to the partial distinguishability model where $x:=1-\epsilon$ [23]. The assumption of identical mutual partial distinguishability is demonstrated to be appropriate in multiphoton experiments with imperfect sources of sufficient quality [58].. A combination of an optical transformation plus partial measurement which projects onto a single photon state $\rho(\epsilon^{\prime})$ results in photon distillation when $\epsilon^{\prime}<\epsilon$.
@@ III Fourier transform-based distillation
Fig. 3 shows the heralding probabilities associated with each family of distillation outcomes. For asymptotically low initial error, we find a total herald probability of $p=\frac{33}{125}+\mathcal{O}(\epsilon)$, divided over many heralding outcomes. This breaks the previously conjectured $p=\frac{1}{N}+\mathcal{O}(\epsilon)$ scaling [17] coming from a single outcome. We strengthen our observations at $N=5$ by numerical simulations up to $N=10$. We find error scaling of $\epsilon^{\prime}=\frac{\epsilon}{N}+\mathcal{O}(\epsilon^{2})$, while simultaneously a total herald success probability of $p\geq\frac{1}{4}+\mathcal{O}(\epsilon)$ is observed [38]. A phenomenon which occurs at higher $N$ which is not observed at $N=5$ is that there are allowed outcomes which nonetheless have zero output probability - a consequence of the fact that Eq. 6 is a necessary but not sufficient condition.
@@ IV Theoretical analysis
We will first focus on the error reduction. In the orthogonal bad-bit model, we have a classical mixture of states containing various error contributions, and error reduction is most easily understood in a Bayesian context: the posterior probability of our state containing zero errors after a partial measurement is given by:
$P({\rho_{\text{i}}}^{\otimes N}|M)=\frac{q_{\text{i}}}{q(\epsilon)}\cdot P({\rho_{\text{i}}}^{\otimes N}).$ (7)
where $q(\epsilon)$ is the probability of obtaining that measurement outcome as a function of the error probability, and $q_{i}$ is the probability of obtaining that measurement outcome for fully indistinguishable photons. This implies that if we observe a partial measurement outcome which is more likely to occur when the state contains zero errors, then we increase the posterior probability that the parts of our state which remain correspond to the part that contains zero errors. We are therefore looking for partial measurements where the probability of observing that outcome $q(\epsilon)$ is a strongly decreasing function of the error probability.
@@
where $\lambda$ is the conditional probability of measuring a given herald measurement outcome given that one of the photons is distinguishable, we find $\lambda=q_{\text{i}}$/$N$. Using a similar approach as Ref. [17], we find for the new error $\epsilon^{\prime}$
@@
We remark that our conjecture sheds some light on why Fourier matrices repeatedly emerge as such useful tools for measuring and controlling partial distinguishability: Eq. 10 also holds trivially for outcomes where all photons emerge from the same mode, in any interferometer, whereas we find that for general outcomes Haar-random interferometers, it does not hold. Outcomes where all photons emerge from the same mode exhibit bosonic bunching, which is a multiphoton constructive interference phenomenon that strongly modifies the output distribution of indistinguishable particles. However, in any interferometer, bunched outcomes only occur with a small probability, which moreover decreases as a function of system size [39]. In a Fourier matrix, in contrast, in the low error regime, our results show that all outcomes behave as if they corresponded to bunched outcomes. We leave this point for future study.
@@
A second approach is concerned only with the limit of large $N$. We can make a few observations. First, we note that for any $N$, the symmetry properties of Fourier matrices mean that they obey the condition $\forall(l_{i}=m_{i}+1\mod N)\rightarrow p_{l}=p_{m}$, where $l$ and $m$ are mode assignment lists. In other words, if all elements of two mode assignment lists differ by one, then the probabilities are equal. Secondly, the substitution $U_{i,j}=\bar{U}_{i,j}$, where the overbar represents the complex conjugate, results in a relabeling of the modes while not altering the permanent of submatrices of $U$. These two relations define equivalence classes in the space of outputs. Third, in the limit of large $N$, almost every allowed belongs to an equivalence class that also contains allowed outcomes which correspond to photon distillation, i.e. which have exactly one photon in the mode of interest. These considerations suggest that in the limit of large $N$, imposing the restriction that there is a single photon in mode $N$ does not substantially modify the statistics of the subset of outcomes that arises as a result of imposing that condition. In other words, a sample of outcomes with this condition will have the same average value as a sample without this condition. In that case, a simple counting argument suffices to compute the limiting value of the success probability: there are $\binom{2N-1}{N}$ measurements in general of which a fraction $\frac{1}{N}$ satisfies the complementary of Eq. 6 [28]. Furthermore, there are $\binom{2N-3}{N-1}$ measurements in which there is exactly 1 photon in output mode $N$, of which a fraction $\frac{1}{N}$ satisfies the complementary of Eq. 6 too. We find a total herald probability
@@ V Discussion
as an upper bound for the scaling coefficient for the scaling coefficient.
@@ VI Implementation in photonic quantum computers
Figure 5: Photon distillation resource allocation tradeoff. This figure shows the tradeoff between allocating photon resources for either error mitigation (by means of photon distillation) or enhancing the quantum error-correction code distance. The use of $N$-photon distillation schemes is advantageous when the initial error rate fraction $p_{\text{error}}/p_{\text{th}}$ exceeds the black curve. Generally, the allocation of photon resources to photon distillation is effective when $p_{\text{error}}/p_{\text{th}}>0.63$.
@@
Establishing this correspondence allows us to make a direct comparison: the error threshold for photon distillation scales as $\epsilon^{*}\approx 1/N^{0.61}$, giving results of $\epsilon\simeq 0.2-0.35$, for the regime where Fig. 5 suggests photon distillation is most resource-efficient. In contrast, thresholds for QEC are $p_{\text{th}}=\mathcal{O}(10^{-4}–10^{-1})$ [4, 45], with FBQC having a threshold of up to 1.07% Pauli error [4].
@@
Finally, we return to the question of whether the use of the small-error limit is justified in these calculations. For our result below threshold, it is clear that this is the case if we, for example, assume the FBQC threshold of 1.07% error. For the results above the QEC threshold, as noted above, even with a pessimistic assumption about $N_{\textrm{gate}}$, we can show that the theshold of the combined scheme exceeds that of bare QEC. However, a precise calculation of the threshold of a combined photon distillation plus QEC scheme is beyond the present work, as this requires both a better understanding of the high-error regime as well as a better understanding of the relationship between partial distinguishability errors and logical errors in photonic quantum computation. We defer the generalization of including other errors, such as photon loss, for future work.
@@ Appendix B Simulation methods
Figure 7: Distillation simulation setup. $N$ partial distinguishable photons are transformed by a $N$-mode Fourier transform $U$ (Eq. 5) into a multiphoton state. The partial state in the first $N-1$ ancillary modes is measured. A successful measurement heralds the presence of a single photon state with reduced error $\epsilon^{\prime}<\epsilon$ (Eq. 1) in mode $N$. The new error $\epsilon^{\prime}$ is extracted by first evaluating the Hong-Ou-Mandel visibility of a distilled photon and a non-distilled photon by interference in a 50:50 beamsplitter. One unitary transformation matrix simultaneously describes all transformations to simulate the interference of $N+1$ partially distinguishable photons.
@@ Appendix C Multi-photon interference coefficients
$c_{j}:=\frac{1}{\mu_{i}}\sum_{\sigma^{j}}\text{perm}[M\circ M_{\sigma^{j}}^{*}].$ (27)
Here, $\mu_{i}$ is a multiplicity correction $\mu(s)=\prod_{i}(s_{i})!$ for outputs with more than one photon in the same mode, $M$ is an $N$-by-$N$ submatrix of unitary transformation matrix $U$ (where rows and columns of $M$ correspond to input and output modes, respectively), $\sigma^{j}$ corresponds to all row permutations $\sigma$ with $N-j$ fixed-point elements, $\circ$ denotes the element-wise matrix product and $M_{\sigma^{j}}^{*}$ is the complex conjugate of the matrix $M$ with rows permuted according to $\sigma^{j}$ (for detail see [23]).
@@
Finally, we summarize the results of our numerical studies for $N=3$ through $N=10$ in Tab. 1. For each allowed outcome with exactly one photon in mode $N$, we calculated the multiphoton coefficients. For each of these allowed outcomes, we checked our conjecture (Eq. 10) by calculating the difference

### v4

@@ I Introduction
The key resource for all linear-optical quantum computing schemes is photon indistinguishability [8, 9, 10], i.e. the degree to which the photons in the quantum computer are identical in the degrees of freedom of those photons which are not being used to encode information (e.g. polarization, time, and frequency [11]). Photonic indistinguishability is essential because while linearity forbids an explicit photon-photon interaction, it allows for multi-particle quantum interference, which is sufficient for universal quantum computation when combined with measurements and fast feedback. The archetypal example of such interference is the Hong-Ou-Mandel effect [12]. Since photon indistinguishability governs the degree to which multi-particle interference occurs, it is the crucial resource for photonic quantum computing in its modern implementations [13, 14, 15, 16], and imperfect indistinguishability gives rise to logical errors in the computation.
@@
In this work, we present a photon distillation scheme with significantly reduced resource costs, rendering the protocol feasible for implementation in photonic quantum computers. Our photon distillation scheme requires no additional technological developments beyond those already needed for fault-tolerant photonic quantum computing architectures. In practice, the resource costs of our quantum error mitigation protocol is expected to be only a small fraction of the total resource costs of the architecture [19]. Our scheme is based on multiphoton interference in Fourier matrices. In contrast to the state of the art [18], our scheme can achieve arbitrary reduction of errors in a single round of the protocol, thereby eliminating the need for concatenation of multiple distillation steps or intermediate fast feedback. Up to a conjecture in the theory of matrix permanents, we show that the resource costs as expressed in the overall number of photons required for the protocol are linear in the required error reduction, and the success probability of the protocol tends to a constant as more photons are used. These two facts render our protocol highly scaleable: to achieve an asymptotic reduction in indistinguishability error by at least a factor of 100 required 59,000 photons with the state-of-the-art scheme, whereas our new scheme only requires 400 photons. We show that our scheme is asymptotically optimal in the degree of error reduction per photon consumed.
@@ II Interference of partially distinguishable photons
We begin by introducing a model for partially distinguishable photons. The orthogonal bad-bit model describes single photon states produced by imperfect sources [17]. The internal error modes do not contribute to multiphoton interference because of their mutual orthogonality, thus allowing us to expand the imperfect single photon states $\rho(\epsilon)$ into equivalent sums of fully indistinguishable $\rho_{\text{i}}=\ket{\psi_{0}}\bra{\psi_{0}}$ | and fully distinguishable $\rho_{\text{d}}$ photons [25]
@@
where $\epsilon$ is the parameter used in this work for the partial distinguishability error  [^1] The orthogonal bad-bit model is equivalent to the partial distinguishability model where $x:=1-\epsilon$ [24]. The assumption of identical mutual partial distinguishability is demonstrated to be appropriate in multiphoton experiments with imperfect sources of sufficient quality [61].. A combination of an optical transformation plus partial measurement which projects onto a single photon state $\rho(\epsilon^{\prime})$ results in photon distillation when $\epsilon^{\prime}<\epsilon$  [^2] We note that for a distilled photon, $\rho_{\text{d}}$ is no longer a pure state but instead a maximally mixed state over the error modes of the initial resource photons. However, the uniqueness of these error modes is preserved in the mixed error state of the distilled photon. Consequently, the precise nature of $\rho_{\text{d}}$ is irrelevant; for subsequent multi-photon interference, it is sufficient to treat $\rho_{\text{d}}$ as equivalent to a pure, fully distinguishable photon state..
@@ III Fourier transform-based distillation
Fig. 3 shows the heralding probabilities associated with each family of distillation outcomes. For asymptotically low initial error, we find a total herald probability of $p=\frac{33}{125}+\mathcal{O}(\epsilon)$, divided over many heralding outcomes. This breaks the previously conjectured $p=\frac{1}{N}+\mathcal{O}(\epsilon)$ scaling [18] coming from a single outcome. We strengthen our observations at $N=5$ by numerical simulations up to $N=10$. We find error scaling of $\epsilon^{\prime}=\frac{\epsilon}{N}+\mathcal{O}(\epsilon^{2})$, while simultaneously a total herald success probability of $p\geq\frac{1}{4}+\mathcal{O}(\epsilon)$ is observed [40]. A phenomenon which occurs at higher $N$ which is not observed at $N=5$ is that there are allowed outcomes which nonetheless have zero output probability - a consequence of the fact that Eq. 6 is a necessary but not sufficient condition. Ref. [26] explicitly demonstrates that for $N$ not equal to prime powers, suppressed allowed outcomes still persist, whereas for $N$ equal to prime powers, Eq. 6 provides a necessary and sufficient condition.
@@ IV Theoretical analysis
We will first focus on the error reduction. In the orthogonal bad-bit model, we have a classical mixture of states containing various error contributions, and the photon distillation circuit transforms this state to:
$\begin{split}\sigma(\epsilon)&=(1-N\epsilon)\mathcal{U}{\rho_{\text{i}}}^{\otimes N}\mathcal{U}^{\dagger}\\ &+\epsilon\sum_{k=1}^{N}\mathcal{U}\rho_{\text{k}}\mathcal{U}^{\dagger}+\mathcal{O}(\epsilon^{2}),\end{split}$ (7)
where we used $\mathcal{U}:=\mathcal{U}(U)$ and $\rho_{\text{k}}:={\rho_{\text{i}}}^{\otimes k-1}\otimes{\rho_{\text{d}}}\otimes{\rho_{\text{i}}}^{\otimes N-k}$ for convenience. The probability of our state containing zero errors after a measurement $M$ is given by:
$\begin{split}&P({\rho_{\text{i}}}^{\otimes N}|M):=\\ &(1-N\epsilon)\langle n_{1},\dots,n_{N-1}1|\mathcal{U}{\rho_{\text{i}}}^{\otimes N}\mathcal{U}^{\dagger}|n_{1},\dots,n_{N-1}1\rangle,\end{split}$ (8)
which signifies that there is 1 photon in mode $N$, as required for photon distillation. Since this distilled photon originates from the zero-error state, it is, by definition, projected into the indistinguishable photon state: $P(\rho_{\text{i}}|\rho_{\text{i}}^{\otimes N})=1$. Similarly, the probability of our state containing one error in input mode $k$ after the same measurement is given by:
$P(\rho_{\text{k}}|M):=\epsilon\langle n_{1},\dots,n_{N-1}1|\mathcal{U}\rho_{\text{k}}\mathcal{U}^{\dagger}|n_{1},\dots,n_{N-1}1\rangle.$ (9)
As the photon distillation transform is symmetric, there is no bias towards a specific output mode for the distinguishable photon. Consequently, there is a $1/N$ probability that the distilled photon will be projected into an error mode, implying that the probability of it being projected into the indistinguishable photon state is $P(\rho_{\text{i}}|\rho_{\text{k}})=(N-1)/N$.
The error reduction is most easily understood in a Bayesian context: the posterior probability of our state containing zero errors after a measurement $M$ is given by:
$P({\rho_{\text{i}}}^{\otimes N}|M)=\frac{q_{\text{i}}}{q(\epsilon)}\cdot P({\rho_{\text{i}}}^{\otimes N}),$ (10)
where $q(\epsilon)$ is the probability of obtaining that measurement outcome as a function of the error probability, and $q_{i}$ is the probability of obtaining that measurement outcome for fully indistinguishable photons. Similarly, the probability of our state containing one error after a measurement $M$ is given by:
$\sum_{k=1}^{N}P(\rho_{\text{k}}|M)=\frac{\lambda}{q(\epsilon)}\cdot\sum_{k=1}^{N}P(\rho_{\text{k}}),$ (11)
where $\lambda$ is the probability of obtaining the same measurement outcome given that one of the photons is distinguishable. This implies that if we observe a partial measurement outcome which is more likely to occur when the state contains zero errors, i.e. $q_{\text{i}}>\lambda$, then we increase the posterior probability that the parts of our state which remain correspond to the part that contains zero errors. We are therefore looking for partial measurements where the probability of observing that outcome, $q(\epsilon)$ is a strongly decreasing function of the error probability. Finally, the new error is defined by the conditional probability of observing an indistinguishable photon given the distilled photon state, $P(\rho_{\text{i}}|\rho(\epsilon^{\prime})):=1-\epsilon^{\prime}$, which is found by evaluating:
$\begin{split}P(\rho_{\text{i}}|\rho(\epsilon^{\prime}))&=P({\rho_{\text{i}}}^{\otimes N}|M)\cdot P(\rho_{\text{i}}|\rho_{\text{i}}^{\otimes N})\\ &+\sum_{k=1}^{N}P(\rho_{\text{k}}|M)\cdot P(\rho_{\text{i}}|\rho_{\text{k}}).\end{split}$ (12)
@@
we find $\lambda=q_{\text{i}}$/$N$. Following a similar approach as Ref. [18] by using Eq. 12, we find for the new error $\epsilon^{\prime}$
@@
We remark that our conjecture sheds some light on why Fourier matrices repeatedly emerge as such useful tools for measuring and controlling partial distinguishability: Eq. 15 also holds for outcomes where all photons emerge from the same mode, in any interferometer, whereas we find that for general outcomes in Haar-random interferometers, it does not hold (App. C). Outcomes where all photons emerge from the same mode exhibit bosonic bunching, which is a multiphoton constructive interference phenomenon that strongly modifies the output distribution of indistinguishable particles. However, in any interferometer, single-mode bunched outcomes only occur with a small probability, which moreover decreases as a function of system size [41]. Counterintuitively, the probability of multimode bunched outcomes is not always maximized for indistinguishable photons [42]. In a Fourier matrix, in contrast, in the low error regime, our results show that all outcomes behave as if they corresponded to single-mode bunched outcomes. We leave this point for future study.
@@
A second approach is concerned only with the limit of large $N$. We can make a few observations. First, we note that for any $N$, the symmetry properties of Fourier matrices mean that they obey the condition $\forall(l_{i}=m_{i}+1\mod N)\rightarrow p_{l}=p_{m}$, where $l$ and $m$ are mode assignment lists. In other words, if all elements of two mode assignment lists differ by one, then the probabilities are equal. Secondly, the substitution $U_{i,j}=\bar{U}_{i,j}$, where the overbar represents the complex conjugate, results in a relabeling of the modes while not altering the permanent of submatrices of $U$. These two relations define equivalence classes in the space of outputs. Third, in the limit of large $N$, almost every allowed outcome belongs to an equivalence class that also contains allowed outcomes which correspond to photon distillation, i.e. which have exactly one photon in the mode of interest. These considerations suggest that in the limit of large $N$, imposing the restriction that there is a single photon in mode $N$ does not substantially modify the statistics of the subset of outcomes that arises as a result of imposing that condition. In other words, a sample of outcomes with this condition will have the same average value as a sample without this condition. In that case, a simple counting argument suffices to compute the limiting value of the success probability: there are $\binom{2N-1}{N}$ measurements in general of which a fraction $\frac{1}{N}$ satisfies the complementary of Eq. 6 [30]. Furthermore, there are $\binom{2N-3}{N-1}$ measurements in which there is exactly 1 photon in output mode $N$, of which a fraction $\frac{1}{N}$ satisfies the complementary of Eq. 6 too. We find a total herald probability
@@ V Discussion
as an upper bound for the scaling coefficient.
@@ VI Implementation in photonic quantum computers
Figure 5: Photon distillation resource allocation trade-off. This figure shows the trade-off between allocating photon resources for either error mitigation (by means of photon distillation) or enhancing the quantum error-correction code distance. The use of $N$-photon distillation schemes is advantageous when the initial error rate fraction $p_{\text{error}}/p_{\text{th}}$ exceeds the black curve. Generally, the allocation of photon resources to photon distillation is effective when $p_{\text{error}}/p_{\text{th}}>0.63$.
@@
Establishing this correspondence allows us to make a direct comparison: the error threshold for photon distillation scales as $\epsilon^{*}\approx 1/N^{0.61}$ [26], giving results of $\epsilon\simeq 0.2-0.35$, for the regime where Fig. 5 suggests photon distillation is most resource-efficient. In contrast, thresholds for QEC are $p_{\text{th}}=\mathcal{O}(10^{-4}–10^{-1})$ [4, 48], with FBQC having a threshold of up to 1.07% Pauli error [4].
@@
Finally, we return to the question of whether the use of the small-error limit is justified in these calculations. For our result below threshold, it is clear that this is the case if we, for example, assume the FBQC threshold of 1.07% error. For the results above the QEC threshold, as noted above, even with a pessimistic assumption about $N_{\textrm{gate}}$, we can show that the theshold of the combined scheme exceeds that of bare QEC. However, a precise calculation of the threshold of a combined photon distillation plus QEC scheme is beyond the present work, as this requires both a better understanding of the high-error regime as well as a better understanding of the relationship between partial distinguishability errors and logical errors in photonic quantum computation. Furthermore, each additional component added to photonic quantum computing architectures introduces new losses, and photon distillation gates are no exception. Therefore, photon distillation gates inherently reduce indistinguishability errors, but at the expense of increased loss errors. In App. E, we discuss the error scaling of realistic photon distillation gates in the presence of loss. We defer the full generalization of including other error types for future work.
@@ Appendix B Simulation methods
Figure 7: Distillation simulation setup. $N$ partial distinguishable photons are transformed by a $N$-mode Fourier transform $U$ (Eq. 5) into a multiphoton state. The partial state in the first $N-1$ auxiliary modes is measured. A successful measurement heralds the presence of a single photon state with reduced error $\epsilon^{\prime}<\epsilon$ (Eq. 1) in mode $N$. The new error $\epsilon^{\prime}$ is extracted by first evaluating the Hong-Ou-Mandel visibility of a distilled photon and a non-distilled photon by interference in a 50:50 beamsplitter. One unitary transformation matrix simultaneously describes all transformations to simulate the interference of $N+1$ partially distinguishable photons.
@@ Appendix C Multi-photon interference coefficients
$c_{j}:=\frac{1}{\mu_{i}}\sum_{\sigma^{j}}\text{perm}[M\circ\bar{M}_{\sigma^{j}}].$ (32)
Here, $\mu_{i}$ is a multiplicity correction $\mu(s)=\prod_{i}(s_{i})!$ for outputs with more than one photon in the same mode, $M$ is an $N$-by-$N$ submatrix of unitary transformation matrix $U$ (where rows and columns of $M$ correspond to input and output modes, respectively), $\sigma^{j}$ corresponds to all row permutations $\sigma$ with $N-j$ fixed-point elements, $\circ$ denotes the element-wise matrix product and $\bar{M}_{\sigma^{j}}$ is the complex conjugate of the matrix $M$ with rows permuted according to $\sigma^{j}$ (for detail see [24]).
@@
We summarize the results of our numerical studies for $N=3$ through $N=10$ in Tab. 1. For each allowed outcome with exactly one photon in mode $N$, we calculated the multiphoton coefficients. For each of these allowed outcomes, we checked our conjecture (Eq. 15) by calculating the difference
@@
Finally, we prove Eq. 15 for single-mode bunched outcomes. First, we combine Eq. 14 and 17 to find $\sum_{j=0}^{N}c_{j}j=(q_{\text{i}}-\lambda)N$. Second, we note that the single-mode bunching probability $p_{\text{single}}=\text{perm}[S]c_{0}$ is determined by the indistinguishability matrix $S$ [42]. For fully indistinguishable photons, $\text{perm}[S]=N!$ which results in $q_{\text{i}}=N!c_{0}$. However, if there is one fully distinguishable photon, $\text{perm}[S]=(N-1)!$ because there exists a row and a column in $S$ for which all the elements, except for the diagonal, are zero. Therefore, we find $\lambda=(N-1)!c_{0}$. Combining the above results, we derive $\sum_{j=0}^{N}c_{j}j=(N-1)N!c_{0}=(N-1)\sum_{j=0}^{N}c_{j}$, which completes our proof.